# Pure-Pass: Fine-Grained, Adaptive Masking for Dynamic Token-Mixing Routing in Lightweight Image Super-Resolution

### v1

@@ Abstract
Image Super-Resolution (SR) aims to reconstruct high-resolution images from low-resolution counterparts, but the computational complexity of deep learning-based methods often hinders practical deployment. CAMixer is the pioneering work to integrate the advantages of existing lightweight SR methods and proposes a content-aware mixer to route token mixers of varied complexities according to the difficulty of content recovery. However, several limitations remain, such as poor adaptability, coarse-grained masking and spatial inflexibility, among others. We propose Pure-Pass ($P$), a pixel-level masking mechanism that identifies pure pixels and exempts them from expensive computations. $P$ utilizes fixed color center points to classify pixels into distinct categories, enabling fine-grained, spatially flexible masking while maintaining adaptive flexibility. Integrated into the state-of-the-art ATD-light model, $P$-ATD-light achieves superior SR performance with minimal overhead, outperforming CAMixer-ATD-light in reconstruction quality and parameter efficiency when saving a similar amount of computation.
@@ Introduction
To mitigate the aforementioned limitations, we propose a novel approach termed Pure-Pass ($P$). In contrast to CAMixer’s window-level masking strategy, our method introduces a more fine-grained pixel-level masking mechanism that selectively identifies pure pixels within homogeneous regions using a novel method based on fixed color centers. These identified pixels are then exempted from expensive computations in the AC-MSA module of ATD-light. To ensure fidelity, a zero-cost compensation mechanism re-integrates information from the parallel SW-MSA branch for these bypassed pixels.
The design of Pure-Pass directly addresses the identified flaws. It offers: (1) Adaptive Flexibility by dynamically determining the amount of computation to save based on image content. (2) Fine-Grained Masking by decoupled window processing from Window Attention for adaptive texture analysis. (3) Spatially Flexible Masking through a cross-shift fusion strategy. (4) Compatibility-Preserving Optimization for SW-MSA by targeting a different module (AC-MSA) with pixel-level masking. (5) Negligible Overhead in parameters and computation. (These are detailed in the Method section). Consequently, our $P$-ATD-light significantly outperforms CAMixer-ATD-light in both reconstruction quality and parameter efficiency, while achieving comparable or even greater computational savings.
@@
We introduce a novel Pure-Pass ($P$) method to efficiently identify texture-deficient regions in input images. Our method generates pixel-level masks through a three-stage process, as illustrated in Figure 2.
@@
To enhance the spatial flexibility of our $P$ method, we introduce a Cross-Shift Mask Fusion strategy. This approach combines masks generated from shifted windows to avoid the dilemma shown in Figure 1.
@@
$P$-ATD-light: Pure-Pass Accelerated ATD-light
@@
    Compatibility-Preserving Optimization for SW-MSA: Our $P$ method generates pixel-level masks, enabling computational optimization for non-Window-Attention modules. Within the ATD-light architecture, which incorporates three parallel attention modules, the computational load is primarily dominated by AC-MSA and SW-MSA operations. Rather than compromising the window-shifting mechanism of SW-MSA through simplified complexity reduction, we implement a targeted optimization strategy that exclusively enhances the computational efficiency of AC-MSA. This methodology preserves the complete functionality of SW-MSA while achieving substantial computational savings.
@@
    Negligible Parametric and Computational Overhead The incorporation of $P$ per se increases the parameter count by less than 1K and the FLOPs by under 0.001G, which is negligible accounting for ATD-light’s total parameter budget and computational load as shown in Table 2.
@@
In the Introduction section, we have highlighted the advantages of the masks generated by our proposed $P$ method compared to those produced by CAMixer. Here, we present empirical evidence through comparative visualizations to substantiate these claims.
The first row of Figure 3 demonstrates the superior performance of $P$-generated masks. As illustrated in Figure 3(b), the CAMixer-derived mask fails to identify homogeneous regions such as the uniform blue sky areas near the building edges and the black building window regions. In contrast, Figure 3(c) shows that mask generated by $P$ is more fine-grained and spatially flexible, enabling the identification of a greater number of computationally efficient regions for SR and thus saving more computational overhead.
The second row of Figure 3 further illustrates the limitations of CAMixer’s fixed masking ratio, which results in unreasonable masking of textured regions such as vegetation and temple eaves (Figure 3(e)). Our $P$ method, however, exhibits dynamic adaptability by selectively masking fewer areas (Figure 3(f)), thereby preserving computational resources for regions that truly benefit from SR processing.
@@
Table 1: Quantitative comparison of super-resolution performance across different datasets and scale factors. The proposed $P$-ATD-light maintains competitive performance with the original ATD-light while significantly outperforming both CAMixer variants, particularly on the challenging Urban100 dataset.
@@
Table 2: Model complexity comparison showing parameter counts and computational requirements. $P$-ATD-light achieves superior efficiency through its adaptive computation mechanism, maintaining the parameter count of ATD-light while reducing FLOPs by 9% on average and up to 21% in optimal cases compared to the original ATD-light.
@@
Tables 1 and 2 present comprehensive comparisons between our proposed $P$-ATD-light and two baseline models: the original ATD-light and CAMixer-ATD-light variants. In CAMixer-ATD-light-ratioratio, the ratioratio hyperparameter determines the proportion of windows performing full self-attention computations, with higher values indicating more computationally intensive processing. Notably, CAMixer-ATD-light-0.5 exhibits inferior performance compared to CAMixer-ATD-light-0.8, demonstrating the trade-off between computational savings and model accuracy.
As evidenced in table 1, $P$-ATD-light achieves comparable performance to ATD-light across all benchmark datasets, while both CAMixer-ATD-light variants show significant performance degradation. This performance gap is particularly pronounced on the Urban100 dataset, where CAMixer-ATD-light-0.8 and CAMixer-ATD-light-0.5 exhibit PSNR drops of 0.32 dB and 0.46 dB respectively compared to our method on $\times 2$ scale.
The efficiency advantages of our approach are detailed in table 2. $P$-ATD-light maintains the parameter efficiency of the original ATD-light (769K parameters), while CAMixer variants require 9% more parameters (838K). When it comes to the FLOPs, as $P$ incorporates an adaptive mechanism that autonomously determines the proportion of regions requiring less intensive computation based on the inherent characteristics of the input images, the percentage of pure pixels for $P$-ATD-light is averaged on Manga109 x2 dataset when calculating FLOPs. Table 2 shows that the average FLOPs of 79.30G for $P$-ATD-light is smaller than the FLOPs of 83.67G for CAMixer-ATD-light-0.8. To demonstrate the potential of the adaptive flexibility feature of $P$, we also provide the maximal percentage of pure pixels of 89.5% when processing an image from B100. And under this percentage, the FLOPs decrease significantly to 69.09G, outperforming even the more aggressive CAMixer-ATD-light-0.5 (73.48G). Visual examples for qualitative comparisons are provided in the supplementary material.
@@
As shown in table 3, our CSMF method increases the percentage of pure pixels from 31.71% to 38.89%, saving 22.66% more FLOPs compared to $P$-ATD-light without CSMF (7.85G v.s. 6.4G). Notably, these computational savings are achieved without compromising model performance, as evidenced by the nearly identical PSNR scores on both Urban100 and Manga109 datasets. This compelling result validates the effectiveness of our CSMF strategy in optimizing computational resources while maintaining reconstruction quality.
@@ Conclusion
In this work, we address key limitations of the existing implementation of Dynamic Token-Mixing Routing strategy for lightweight image super-resolution, such as poor adaptability, coarse-grained masking, and spatial inflexibility. By introducing Pure-Pass ($P$), a pixel-level masking mechanism, we enable fine-grained, spatially flexible masking that exempts pure pixels from costly computations while preserving adaptive flexibility. $P$ leverages fixed color center points to categorize pixels efficiently, ensuring minimal computational overhead. When integrated into the state-of-the-art ATD-light model, $P$-ATD-light demonstrates superior SR performance, surpassing CAMixer-ATD-light in both reconstruction quality and parameter efficiency while maintaining comparable computational savings. Our approach advances the practical deployment of efficient SR models by balancing performance and complexity.

### v2

@@ Abstract
Image Super-Resolution (SR) aims to reconstruct high-resolution images from low-resolution counterparts, but the computational complexity of deep learning-based methods often hinders practical deployment. CAMixer is the pioneering work to integrate the advantages of existing lightweight SR methods and proposes a content-aware mixer to route token mixers of varied complexities according to the difficulty of content recovery. However, several limitations remain, such as poor adaptability, coarse-grained masking and spatial inflexibility, among others. We propose Pure-Pass (PP), a pixel-level masking mechanism that identifies pure pixels and exempts them from expensive computations. PP utilizes fixed color center points to classify pixels into distinct categories, enabling fine-grained, spatially flexible masking while maintaining adaptive flexibility. Integrated into the state-of-the-art ATD-light model, PP-ATD-light achieves superior SR performance with minimal overhead, outperforming CAMixer-ATD-light in reconstruction quality and parameter efficiency when saving a similar amount of computation.
@@ Introduction
To mitigate the aforementioned limitations, we propose a novel approach termed Pure-Pass (PP). In contrast to CAMixer’s window-level masking strategy, our method introduces a more fine-grained pixel-level masking mechanism that selectively identifies pure pixels within homogeneous regions using a novel method based on fixed color centers. These identified pixels are then exempted from expensive computations in the AC-MSA module of ATD-light. To ensure fidelity, a zero-cost compensation mechanism re-integrates information from the parallel SW-MSA branch for these bypassed pixels.
The design of Pure-Pass directly addresses the identified flaws. It offers: (1) Adaptive Flexibility by dynamically determining the amount of computation to save based on image content. (2) Fine-Grained Masking by decoupled window processing from Window Attention for adaptive texture analysis. (3) Spatially Flexible Masking through a cross-shift fusion strategy. (4) Compatibility-Preserving Optimization for SW-MSA by targeting a different module (AC-MSA) with pixel-level masking. (5) Negligible Overhead in parameters and computation. (These are detailed in the Method section). Consequently, our PP-ATD-light significantly outperforms CAMixer-ATD-light in both reconstruction quality and parameter efficiency, while achieving comparable or even greater computational savings.
@@
We introduce a novel Pure-Pass (PP) method to efficiently identify texture-deficient regions in input images. Our method generates pixel-level masks through a three-stage process, as illustrated in Figure 2.
@@
To enhance the spatial flexibility of our PP method, we introduce a Cross-Shift Mask Fusion strategy. This approach combines masks generated from shifted windows to avoid the dilemma shown in Figure 1.
@@
PP-ATD-light: Pure-Pass Accelerated ATD-light
@@
    Compatibility-Preserving Optimization for SW-MSA: Our PP method generates pixel-level masks, enabling computational optimization for non-Window-Attention modules. Within the ATD-light architecture, which incorporates three parallel attention modules, the computational load is primarily dominated by AC-MSA and SW-MSA operations. Rather than compromising the window-shifting mechanism of SW-MSA through simplified complexity reduction, we implement a targeted optimization strategy that exclusively enhances the computational efficiency of AC-MSA. This methodology preserves the complete functionality of SW-MSA while achieving substantial computational savings.
@@
    Negligible Parametric and Computational Overhead The incorporation of PP per se increases the parameter count by less than 1K and the FLOPs by under 0.001G, which is negligible accounting for ATD-light’s total parameter budget and computational load as shown in Table 2.
@@
In the Introduction section, we have highlighted the advantages of the masks generated by our proposed PP method compared to those produced by CAMixer. Here, we present empirical evidence through comparative visualizations to substantiate these claims.
The first row of Figure 3 demonstrates the superior performance of PP-generated masks. As illustrated in Figure 3(b), the CAMixer-derived mask fails to identify homogeneous regions such as the uniform blue sky areas near the building edges and the black building window regions. In contrast, Figure 3(c) shows that mask generated by PP is more fine-grained and spatially flexible, enabling the identification of a greater number of computationally efficient regions for SR and thus saving more computational overhead.
The second row of Figure 3 further illustrates the limitations of CAMixer’s fixed masking ratio, which results in unreasonable masking of textured regions such as vegetation and temple eaves (Figure 3(e)). Our PP method, however, exhibits dynamic adaptability by selectively masking fewer areas (Figure 3(f)), thereby preserving computational resources for regions that truly benefit from SR processing.
@@
Table 1: Quantitative comparison of super-resolution performance across different datasets and scale factors. The proposed PP-ATD-light maintains competitive performance with the original ATD-light while significantly outperforming both CAMixer variants, particularly on the challenging Urban100 dataset.
@@
Table 2: Model complexity comparison showing parameter counts and computational requirements. PP-ATD-light achieves superior efficiency through its adaptive computation mechanism, maintaining the parameter count of ATD-light while reducing FLOPs by 9% on average and up to 21% in optimal cases compared to the original ATD-light.
@@
Tables 1 and 2 present comprehensive comparisons between our proposed PP-ATD-light and two baseline models: the original ATD-light and CAMixer-ATD-light variants. In CAMixer-ATD-light-ratioratio, the ratioratio hyperparameter determines the proportion of windows performing full self-attention computations, with higher values indicating more computationally intensive processing. Notably, CAMixer-ATD-light-0.5 exhibits inferior performance compared to CAMixer-ATD-light-0.8, demonstrating the trade-off between computational savings and model accuracy.
As evidenced in table 1, PP-ATD-light achieves comparable performance to ATD-light across all benchmark datasets, while both CAMixer-ATD-light variants show significant performance degradation. This performance gap is particularly pronounced on the Urban100 dataset, where CAMixer-ATD-light-0.8 and CAMixer-ATD-light-0.5 exhibit PSNR drops of 0.32 dB and 0.46 dB respectively compared to our method on $\times 2$ scale.
The efficiency advantages of our approach are detailed in table 2. PP-ATD-light maintains the parameter efficiency of the original ATD-light (769K parameters), while CAMixer variants require 9% more parameters (838K). When it comes to the FLOPs, as PP incorporates an adaptive mechanism that autonomously determines the proportion of regions requiring less intensive computation based on the inherent characteristics of the input images, the percentage of pure pixels for PP-ATD-light is averaged on Manga109 x2 dataset when calculating FLOPs. Table 2 shows that the average FLOPs of 79.30G for PP-ATD-light is smaller than the FLOPs of 83.67G for CAMixer-ATD-light-0.8. To demonstrate the potential of the adaptive flexibility feature of PP, we also provide the maximal percentage of pure pixels of 89.5% when processing an image from B100. And under this percentage, the FLOPs decrease significantly to 69.09G, outperforming even the more aggressive CAMixer-ATD-light-0.5 (73.48G). Visual examples for qualitative comparisons are provided in the supplementary material.
@@
As shown in table 3, our CSMF method increases the percentage of pure pixels from 31.71% to 38.89%, saving 22.66% more FLOPs compared to PP-ATD-light without CSMF (7.85G v.s. 6.4G). Notably, these computational savings are achieved without compromising model performance, as evidenced by the nearly identical PSNR scores on both Urban100 and Manga109 datasets. This compelling result validates the effectiveness of our CSMF strategy in optimizing computational resources while maintaining reconstruction quality.
@@ Conclusion
In this work, we address key limitations of the existing implementation of Dynamic Token-Mixing Routing strategy for lightweight image super-resolution, such as poor adaptability, coarse-grained masking, and spatial inflexibility. By introducing Pure-Pass (PP), a pixel-level masking mechanism, we enable fine-grained, spatially flexible masking that exempts pure pixels from costly computations while preserving adaptive flexibility. PP leverages fixed color center points to categorize pixels efficiently, ensuring minimal computational overhead. When integrated into the state-of-the-art ATD-light model, PP-ATD-light demonstrates superior SR performance, surpassing CAMixer-ATD-light in both reconstruction quality and parameter efficiency while maintaining comparable computational savings. Our approach advances the practical deployment of efficient SR models by balancing performance and complexity.